A thin film of ethyl alcohol (nf = 1.36) is spread on a glass plate and viewed from directly above as the liquid evaporates. At a certain moment a point on the film appears bright red (λ0 = 700 nm). What is the minimum thickness of the alcohol at that instant? [Hint: Compare the index of refraction of the film to that of the media on either side of it. Is there a phase shift due to internal and external reflection?]
Answer #1

The light we see in reflection is the result of interference between light reflected from the air/ethyl alcohol interface and light reflected from the ethyl alcohol/glass interface. If the region is seen to be primarily green, this means that constructive interference occurs between these regions from light with a wavelength of 700 nm. The condition for constructive interference is satisfied when the light which goes through the film and is reflected back up goes through a path length which is an integer multiple of the wavelength (in ethyl alcohol), i.e.
2d = m?/n
where d is the thickness of the film, m is a positive integer, ? = 700 nm and n = 1.36
Plugging in for your values, we find
d = m (700 nm/(2*1.36)) = m * (1904 nm)